OpenStudy (hannahc234):

I will fan and medal. Simplify (6x^-2) ^2(0.5x)^4. Show your work. I have no clue how to answer this question and this is the last one I need help with. I get all the other ones.

OpenStudy (anonymous):

Do you know the first step?

OpenStudy (anonymous):

What I do is start with the Distributive Property of Multiplication. \[(xy)^a=x^ay^a\] That's the formula

OpenStudy (anonymous):

So if we plug that in: \[6^2(x^-2)^2 (0.5x)^4\] Do you know what to do after that?

OpenStudy (hannahc234):

To be honest I have no clue what to do with this expression.

jhonyy9 (jhonyy9):

so x^-2 = 1/x^2 6x^-2 = 6/x^2 (6x^-2)^2 = 36/x^4 (0,5x)^4 = 0,0625x^4 (36/x^4)*(0,0625x^4) = 36*0,0625 = 2,25 x^4 with x^4 simplified and remain there

OpenStudy (anonymous):

What he said. xD ^

OpenStudy (hannahc234):

Okay with my and my dumb self I got the beginning. But the ending of it was confusing. What's the answer? Tbh that sounded so complicated sorry.

OpenStudy (anonymous):

2.25